### v5

# High-performance real-world optical computing trained by in situ gradient-based model-free optimization

Abstract
Optical computing systems provide high-speed and low-energy data processing but face deficiencies in computationally demanding training and simulation-to-reality gaps. We propose a gradient-based model-free optimization (G-MFO) method based on a Monte Carlo gradient estimation algorithm for computationally efficient in situ training of optical computing systems. This approach treats an optical computing system as a black box and back-propagates the loss directly to the optical computing weights’ probability distributions, circumventing the need for a computationally heavy and biased system simulation. Our experiments on diffractive optical computing systems show that G-MFO outperforms hybrid training on the MNIST and FMNIST datasets. Furthermore, we demonstrate image-free and high-speed classification of cells from their marker-free phase maps. Our method’s model-free and high-performance nature, combined with its low demand for computational resources, paves the way for accelerating the transition of optical computing from laboratory demonstrations to practical, real-world applications.
Index Terms:
Computational Optics; Optical Neural Network; Model-free Training;
1 Introduction
Optical computing leverages the properties of light waves to facilitate high-speed data processing while reducing the energy cost [1, 2, 3, 4, 5, 6, 7]. Recent advances in automatic differentiation have enabled in silico training of large-scale optical computing weights, giving rise to the realizations of diffractive neural networks [8, 9], optical reservoir computing [10, 11], and coherent nanophotonic circuits [12].
Problem Statement. Training optical computing systems presents two challenges: an intensive computational process and a performance disparity between simulation and reality when implementing pre-trained weights onto real-world systems [13, 9, 14]. Typically, the optical computing systems are trained in silico using differentiable simulators rooted in the first principle of optics, an approach known as simulator-based training (SBT). While SBT has proven effective within the confines of the simulator, the performance in real systems is largely contingent upon the simulator’s fidelity. Factors such as misalignment and aberration, often omitted in simulations, cause significant performance degradation when optical computing weights trained exclusively within the simulator are applied to real-world systems.
To bridge the reality gap between simulation and experiments, physics-aware training (PAT) and hybrid training (HBT) have been introduced  [15, 16] recently. Both training strategies include conducting the forward pass in the real-world system and back-propagating the loss from the system to its weights through the simulator. These in situ approaches allow the training process to access the optical computing system during the forward pass, which leads to more accurate weight updates than in silico training  [13].
Despite these recent advances, there is a continued reliance on a physics-based simulator during the back-propagation process in current in situ training methods. Such a setting brings three drawbacks: (1) the bias between the simulator and real system prohibits the above training process from achieving optimal results; (2) the in silico simulation requires large memory and computation, limiting the aforementioned methods from in situ training in edge devices with limited computing resources [17]; and (3) the model-based training strategies need high-fidelity images of the input object, which are costly to acquire in terms of instruments, time and storage memory. Hence, developing a memory- and computation-efficient algorithm to train the optical computing system efficiently is an open problem we address in this paper.
Figure 1: Gradient-based model-free optimization based training of the optical computing system. (a) The brown highlights show that our gradient-based model-free training strategy back-propagates the error of training to the distribution parameter $\theta$ and bypasses the reliance on correct differentiable modeling of the optical system $f_{sys}$ and knowledge on input $\{x_{i}\}_{i=1}^{N}$. (b) The blue highlights show that the conventional training of the optical computing system relies on a physics-based simulator $\hat{f}_{sys}$, which substitutes the inaccessible $f_{sys}$ corresponding to the real system. The training process back-propagates the loss through simulator $\hat{f}_{sys}$ to update the weight $w$. This is the basis of SBT and HBT methods.
Proposed Solution. Here, we propose an in situ model-free optimization method that does not require back-propagating errors through the simulator. Instead, we use a score gradient estimation algorithm [18, 19] to solely use the forward outputs from the real system to get gradients for updating the weights of the optical computing system. As shown in Fig. 1, our method treats the optical system as a black box and back-propagates task-specific negative-loss as rewards to the source weight distributions (Fig. 1a). This process only requires knowledge of the weights and forward outputs of the optical computing system, unlike the SBT and HBT methods that require a simulator and the images of the input objects (Fig. 1b).
Contributions. We make the following contributions:
  • We introduce a score-gradient estimation method to train the optical computing system in a model-free manner, providing a computation- and memory-efficient way to mitigate the reality gap to training the real optical computing system (Sec. 3).
  • We validate our method on a diffractive optical computing system. Experimental results show that G-MFO outperforms hybrid training on the commonly used MNIST and FMNIST datasets [9, 15, 16] (Sec. 4.1).
  • We show the G-MFO training process only consumes $\sim 0.1\%$ of the GPU time and memory compared with the HBT in situ method by avoiding the computation-intensive modeling of the wave propagation process (Sec. 4.2).
  • Application scenario: We demonstrate, as a proof-of-concept, that our G-MFO-trained optical computing system’s effectiveness on classifying four types of white blood cells from their phase maps in a marker-free111 Herein, we use ”marker-free” but not ”label-free” to avoid confusion. This is because ”label-free” in biomedical research and high-level vision have different meanings. manner with a testing accuracy of $73.8\%$, making it a promising approach for image-free, marker-free, and high-speed cell analysis (Sec. 4.4).
Open source. Upon acceptance, we will open-source all the code (including simulators, algorithm-related, hardware-related, and baselines) related to our work to facilitate research on training optical computing systems. We hope to promote the reproducibility and follow-ups for this hardware-centric research direction.
Conventional MBO MFO
SBT  [9, 12] H-MBO  [15, 16, 20] L-MBO [21, 22] IMOB [23, 24, 25] FFT  [26, 27] Trial and Error [28, 29] GA  [30] G-MFO (ours)
In situ No Yes Yes Yes Yes Yes Yes Yes
Computation overhead High High High Low Low Low Low Low
Input image free No No No Yes No Yes Yes Yes
In silico training time High High High Low Low Low Low Low
Dimensionality High High High High High Low Low Medium
Intermediate measurement No No No Yes Yes No No No
Gradient Yes Yes Yes Yes No No No Yes
TABLE I: Comparison of strategies on training optical computing systems along the axes of in situ training capability (in situ), in silico computation overhead (computation overhead), the requirement on the recording of the input object to the task (input image free), required in silico training time, the dimensionality of trainable parameters (dimensionality), requirements on intermediate measurement and whether or not use gradient-based update that is more efficient (gradient). *Denotes compatibility with adding nonlinearity to the optical computing system. Simulator-based training (SBT), Hybrid model-based training (H-MBO), Learned model-based training (L-MBO), Intermediate measurement optical backpropagation (IMOB), Forward-forward training (FFT), Genetic algorithm (GA), Gradient-based model-free optimization (G-MFO).
2 Related Work
We discuss the following work related to our contributions.
2.1 In situ training strategies to optimize the optical computing system
The majority of research into optical computing system training has historically relied on in silico training. This process involves carrying out forward and backward propagation calculations on an external computer that simulates the physical system through a digital twin [9, 12]. However, this technique can lead to discrepancies between the simulation and actual reality due to inaccuracies in the physical system’s representation.
In recent years, progress has been made in developing in situ training strategies that use data gathered from real-world optical computing systems to mitigate the reality gap and improve experimental performance. As a first attempt, we categorize previous and our in-situ training methods into two genres: model-free optimization (MBO) and model-based optimization (MFO) methods. We differentiate these two genres by whether they use numerical models of the optical computing system during the parameter updating process. A tabular comparison of the previous and our methods is in Tab. I. Details of the illustrations and comparisons are illustrated below.
MBO methods. Physics-aware training (PAT) [15], hybrid training (HBT) [16] or adaptive training (AT) [20] (H-MBO) are one type of the model-based methods that conduct forward pass in a real system and back-propagate the gradients with the aid of a physics-based simulator. Backpropagation through learned model (L-MBO) is another method that gathers the real data from the optical computing system to train a neural proxy of the optical computing system [21, 22]. With such a learned proxy rather than the physics-based model utilized in PAT/HBT/AT, the backward parameter update process can reduce the bias from the model mismatch between the simulator and the real system. During training, one has to optimize both the task’s and proxy’s parameters to optimize the task performance and realize high fidelity. The task performance for the MBO, especially the L-MBO, is high, while back-propagating with the digital twin also burdens the training process. Moreover, all the MBO approaches require images of input objects during the training, imposing another burden on the overall process.
MFO methods. Unlike the aforementioned MBO methods, MFO methods do not require constructing a digital twin of the optical computing system; they directly use the real system’s output or intermediate measurements to update its parameters. Population-based algorithms such as genetic algorithm (GA) methods [30] use the output from the optical computing system to optimize parameters by simulating the process of natural selection. Still, the genetic algorithm’s performance scales poorly to the dimensionality of parameters [31]. Forward-forward training (FFT) is a greedy multi-layer learning procedure for mortal computation [32] and it has been applied to train optical computing systems reently [26, 27]. However, labels must be added to each input sample during the training and testing. As a result, such training strategies require masking the input, which might be unrealistic in many scenarios, and the testing process of forward-forward training is prolonged. Intermediate measurement optical backpropagation (IMOB) proposed using the adjoint variable method or direct feedback alignment, etc., to calculate gradients [23, 24, 25]. However, such methods require intermediate system measurements for every layer’s output, which might be infeasible or require additional measurement setups.
We propose an MFO training method using Monte Carlo gradient estimation to automatically update the optical computing system parameters. Our work shares similarities with Zhang et al., which uses the genetic algorithm for weight update [30], as our and their methods both generate a batch of sampled parameters during the training process. The difference is that our parameter updating process is more efficient owing to a score gradient estimator rather than their heuristic hill-climbing type algorithm. Moreover, our method does not need specific conditions, such as changing the system or manipulating the input object, which are more or less required in other MBO and MFO methods.
2.2 Application side of optical computing
Optical computing, outperforming electrical computing in parallelism and processing speed, is pivotal for AI progress via optical neural networks (ONN) development [5]. ONNs encode inputs using light’s spatial, temporal, and spectral characteristics to enable applications such as image classification, motion detection, and medical diagnosis [33, 29, 34]. Furthermore, optical computing functions as an advanced neural network platform and directly processes optical signals, facilitating applications like phase reconstruction [35], denoising [36], edge detection [37], and unidirectional imaging [38].
As a secondary contribution, our work showcases the application of optical computing for marker-free cell classification. Particularly relevant to our research is the study by Wang et al. [39], which demonstrates the classification of fluorescent-labeled cells using an optical computing system. In contrast, our approach shows cell-classification based on their phase maps, instead of relying on fluorescent markers. This illustrates that optical computing can facilitate low-storage, high-speed, and marker-free cell analysis.
3 Methodology
In what follows, we mathematically detail the problem setup on training the optical computing system and our solution. We introduce problem formation in Subsec. 3.1 and the conventional solution of using simulator-based training (SBT) in Subsec. 3.2. We then illustrate our solution to the problem, the gradient-based model-free optimization (G-MFO) for training the optical computing system in Subsec. 3.3. Finally, we illustrate the optical computing system and the related simulators in Subsec. 3.4, which we will use to demonstrate the performance of our method.
3.1 Problem setup
We are interested in learning the optimal weight $w\in\mathbb{R}^{H}$ for the optical computing system on a desired task with a training dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ is the size of the dataset, $H$ is the number of trainable parameters in $w$, and $x$ and $y$ denote the input and target of interest, respectively. A function $f_{sys}(\cdot,w)$ maps $x\rightarrow y$ through this optical computing system with $w$. Specifically, in the image classification task based on the diffractive optical computing system we work on, $f_{sys}$ denotes the optical mapping from the input image $x$ to the output label $y$, and $w$ is the optical computing weight in the form of phase value modulation.
During training, one can minimize the cost function $J(w)$ as the mean task-specific loss across the entire training data set $\mathcal{D}$ :
$\operatorname*{arg\,min}_{w}J(w):$ $=\mathbb{E}[\mathcal{L}(\mathcal{D},w)],$ (1a)
$=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w),y_{i}),$ (1b)
where $\mathcal{L}$ is the task-specific loss function, we use cross-entropy loss [40] since we deal with image classification tasks throughout this paper.
We use gradient descent-based search to find the optimal $w$ to minimize the objective function $J(w)$:
$w=w-\alpha\nabla_{w}J(w),$ (2)
where $\nabla_{w}$ represents the gradient operator that collects all the partial derivatives of a function concerning parameters in $w$, and $\alpha$ is the learning rate.
It is straightforward to use the backpropagation method [14] to take the gradient through $f_{sys}$ and finds the gradient $\nabla_{w}J(w)$ as:
$\nabla_{w}J(w)=\frac{1}{N}\sum_{i=1}^{N}\nabla_{w}\mathcal{L}({f_{sys}(x_{i},w)},y_{i}),$ (3)
when we have an accurate and differentiable $f_{sys}$ modeling. However, this is the case for training digital neural networks [14], but not when we train a real-world optical computing system. Thus, this paper’s critical aim is finding an accurate gradient estimation of $\nabla_{w}J(w)$ to update optical computing weight $w$ in a real-world optical system.
Figure 2: System misalignment in a real optical computing system degenerates the performance of the optical computing system trained solely with a physics-based simulator. (a) Testing accuracy drops to $36.4\%$ from $82.2\%$ when having a misaligned rotation angle $\Delta\phi_{I}=0.01\degree$. (b) Testing accuracy drops to $51.0\%$ from $82.2\%$ when the x’-axis misalignment of the optical computing layer $\Delta d_{c}$ is $41.1\>\mu m$. (c) Testing accuracy decreases to $50.9\%$ from $82.2\%$ when the y’-axis misalignment of the output layer $\Delta d_{o}$ is $62.4\>\mu m$.
3.2 In silico simulator-based training (SBT)
Back-propagation through the simulator $\hat{f}_{sys}$. In a real-world optical computing system, since we do not have an exact functional expression of $f_{sys}$, the simulator-based training (SBT) builds a simulator $\hat{f}_{sys}$ as the differentiable approximation of $f_{sys}$ (Fig. 1b). A naive training strategy is substituting $f_{sys}$ in Eq. 3 with the simulator $\hat{f}_{sys}$ and applying in silico training on the simulator:
$\nabla_{w}J(w)=\frac{1}{N}\sum_{i=1}^{N}\nabla_{w}\mathcal{L}({\hat{f}_{sys}(x_{i},w)},y_{i}).$ (4)
The resulting $\nabla_{w}J(w)$ is used in Eq. 2 to update $w$. After training, the optimized $w$ is uploaded to the real optical computing system $f_{sys}$ to test the performance.
Simulation-to-reality gap. The aforementioned simulator-based training is based on backpropagation through the simulator $\hat{f}_{sys}$. The ”sim2real” gap is low (i.e., the gradient $\nabla_{w}J(w)$ is an accurate estimation) when the simulator $\hat{f}_{sys}$ is similar to $f_{sys}$. However, this assumption does not hold in many prototypes of optical computing systems where inadequate modeling and misalignment between the optical elements decay the performance of the SBT during the ”sim2real” transfer. We use a simulator described in Subsec. 3.4 to assess the adverse effect of misalignment on the image classification results. We impose various misalignments to a well-trained ideal optical computing system and measure drops in classification accuracy. For instance, we show in Fig. 2 that slightly laterally misaligning the optical computing layer by $41.1\>\mu m$ reduces the classification accuracy by $31.2\%$.
3.3 In situ G-MFO
Our solution for solving the aforementioned ”sim2real” gap issue in Subsec. 3.2 is in situ learning the optical computing weight $w$ with the gradient-based model-free optimization (G-MFO). In situ learning enables us to access the output of $f_{sys}$ and is feasible on the hardware side because we can use programmable devices such as spatial light modulators to update $w$. The challenging part is designing a training strategy that efficiently uses the output of actual system $f_{sys}$ to construct an unbiased gradient estimator. Here, we use the score gradient estimator to calculate the gradient [41, 42] for the backward update of parameters in $w$ while circumventing the construction of $\hat{f}_{sys}$, a biased and resource-intensive numerical modeling of $f_{sys}$.
Back-propagation through the weights distribution $p$. In our score gradient estimator for model-free optimization, we model optical computing weight $w$ as a random variable that follows a probability distribution: $w\sim p(w|\theta)$ and rewrite the cost function $J(w)$ in Eq. 1b as a probability cost function $J(\theta)$:
$\operatorname*{arg\,min}_{\theta}J(\theta):$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{p(w|\theta)}[\mathcal{L}(f_{sys}(x_{i},w),y_{i})],$ (5a)
$=\frac{1}{N}\sum_{i=1}^{N}\int{p(w|\theta)}\mathcal{L}(f_{sys}(x_{i},w),y_{i})\mathop{}\!\mathrm{d}w,$ (5b)
where the probability distribution $p(w|\theta)$ is continuous in its domain and differentiable concerning its distribution parameter $\theta$. Accordingly, the original goal of optimizing $w$ is reformulated as finding a most likely distribution $p(w|\theta)$ that minimizes the cost function $J(\theta)$ in Eq. 5. Specifically, we model the distribution $p$ as a multivariate normal distribution with $\theta=\{\mu,\sigma^{2}\}$ and $p(w|\theta)=\mathcal{N}(w;\mu,\sigma^{2})$ for optimizing the continuous phase-valued weight to be uploaded onto the SLM in our work.
To update distribution parameter $\theta$ with the gradient descent Eq. 2, we take the gradient of the cost function $J(\theta)$ in Eq. 5:
$\nabla_{\theta}J(\theta)$ $=\nabla_{\theta}\frac{1}{N}\sum_{i=1}^{N}\int{p(w|\theta)}\mathcal{L}(f_{sys}(x_{i},w),y_{i})\mathop{}\!\mathrm{d}w,$ (6a)
$=\frac{1}{N}\sum_{i=1}^{N}\int\mathcal{L}(f_{sys}(x_{i},w),y_{i})\nabla_{\theta}{p(w|\theta)}\mathop{}\!\mathrm{d}w.$ (6b)
Applying the log derivative trick, we have:
$\nabla_{\theta}J(\theta)=\frac{1}{N}\sum_{i=1}^{N}\int p(w|\theta)\mathcal{L}(f_{sys}(x_{i},w),y_{i})\nabla_{\theta}\log{p(w|\theta)}\mathop{}\!\mathrm{d}w.$ (7)
Then we apply Monte Carlo integration to approximate the integral value in Eq. 7 by first drawing $M$ independent samples $\{w_{j}\}_{j=1}^{M}$ from the distribution $p(w|\theta)$ and then computing the average function value evaluated in these samples:
$\nabla_{\theta}J(\theta)$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})\nabla_{\theta}\log{p(w_{j}|\theta)},$ (8a)
$=\frac{1}{M}\sum_{j=1}^{M}[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})]\nabla_{\theta}\log{{p(w_{j}|\theta)}}.$ (8b)
Here we define $r(w_{j})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})$ as the negative reward corresponding to each weight $w_{j}$. We use Equation 8b as the score gradient estimator for G-MFO, where the score function is $\nabla_{\theta}\log{p(w_{j}|\theta)}$, which has been widely used in other areas, such as policy gradient algorithms in reinforcement learning [19] and diffusion models [43].
Variance reduction. The main risk of using the score gradient estimator is the high variance that comes from the Monte Carlo integration step that transits Eq. 7 to Eq. 8a. Such a sampling-based integration step has high variance because different sets of random samples may lead to significantly different integral estimates. We reduce the variance by subtracting the $r(w_{j})$ with baseline value $\bar{r}=\frac{1}{M}\sum_{j=1}^{M}r(w_{j})$ while keeping the bias of gradient unchanged [44]:
$\nabla_{\theta}J(\theta)=\frac{1}{M}\sum_{j=1}^{M}(r(w_{j})-\bar{r})\nabla_{\theta}\log{{p(w_{j}|\theta)}}.$ (9)
Training recipe of G-MFO. We perform the following steps in each training epoch (see Fig. 3): 1. Sample a group of phase-valued optical computing weights $\{w_{j}\}_{j=1}^{M}$ from the distribution $w_{j}\sim p(w|\theta)$ and upload them to the optical computing layer. Upload classification data $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$ to the input layer; 2. Collect the optical computing system’s output $\{f_{sys}(x_{i},w_{j})\}_{i=1,j=1}^{N,M}$ for each pair of input and weight; 3. Update the distribution parameter $\theta$ in silico using Eqs. 2 and 9; 4. Sample a new group of $\{w_{j}\}_{j=1}^{M}$ from the updated distribution $w_{j}\sim p(w|\theta)$, which is ready to be uploaded to the optical computing layer in the next epoch. The algorithm iterates these steps until convergence.
After minimizing the objective function Eq. 5, we export $w_{j}$ with the smallest $r(w_{j})$ as the output weight $w^{\star}$. This performs better than setting $w^{\star}$ as the sampled mean $\mu$. An algorithmic overview of the above training recipe is shown in Algorithm 1.
Algorithm 1 Algorithmic overview of G-MFO.
1:Input: Classification dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$, learning rate $\alpha$, number of sampled weights $M$, optical computing system $f_{sys}$, distribution parameter $\theta=\{\mu,\sigma^{2}\}$, loss function $\mathcal{L}$, epochs $K$.
2:Output: Optimized optical computing weight $w^{\star}$.
3:for $k$ in range $K$ do
4:     Sample $\{w_{j}\}_{j=1}^{M}$ from distribution $p(w|\theta)$.
5:     $\triangleright$ in situ evaluate $\{w_{j}\}_{j=1}^{M}$.
6:     for $j$ in range $M$ do
7:
8:         $r(w_{j})\leftarrow\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i}).$
9:     end for
10:     $\triangleright$ in silico update $\theta$.
11:     Calculate $\nabla_{\theta}J(\theta)$ via Eq. 9.
12:     $\theta\leftarrow\theta-\alpha\nabla_{\theta}J(\theta).$
13:end for
14:$w^{\star}\leftarrow w_{j}$ with the smallest $r(w_{j})$.
Figure 3: An illustration of the training process for the MFO experiment on the real system.
Figure 4: Setups of optical computing systems. (a) In the single-layer optical computing system, we have a light source incident from the top left, passing through the collimator lens (CL), polarizer (P), and the first beam splitter (BS1) and then reflecting from the input layer, which displays phase object by phase delay from the SLM1. The reflected light then passes through BS1 and beam splitter 2 (BS2) and arrives at the optical computing layer, where we use SLM2 to upload the optical processing weights. The reflected light from the optical computing layer then arrives at the output layer, and we calculate the task-specific loss based on the arrival signal. We label the last two paths as $Prop_{IC}$ and $Prop_{CO}$, respectively, in the figure. (b) In the two-layer optical computing system, the light passing through CL and P is reflected by a mirror (M) and directed towards the input layer. The input layer displays binary intensity objects using a DMD. The reflected light then reaches the optical computing layer 1 and, subsequently, optical computing layer 2, where we use SLM1 and SLM2 to upload the optical processing weights. Finally, the light arrives at the output layer, and we obtain the output of the optical computing system. We label the three paths as $Prop_{IC_{1}}$, $Prop_{C_{1}C_{2}}$, $Prop_{C_{2}O}$.
3.4 Experiment and simulation detail of our diffractive optical computing system
3.4.1 Experimental setups of the real single-layer and two-layer optical computing systems $f_{sys-1-layer}$ and $f_{sys-2-layer}$.
We build a single-layer optical computing system and a two-layer optical computing system to validate the effectiveness of the proposed G-MFO-based training strategy (see Fig. 4 and Supplement Fig. S.1).
Single-layer optical computing system: A laser field $u_{laser}$ propagates onto the input layer (SLM1), where the phase-valued object $u_{obj}$ is displayed. The light field reflects from the input layer onto the optical computing layer $u_{w}$ (SLM2). A camera then detects the diffraction light at the output layer. The length $d_{IC}$ and $d_{CO}$ for $Prop_{IC}$ and $Prop_{CO}$ in Fig. 4(a) are $215.1mm$ and $201.6mm$, respectively. As there are multiple components in the system (e.g., SLMs and cameras), it is difficult to manually align them with high accuracy. We developed a digital alignment method based on the homography technique [45] using SLMs to compensate for the alignment errors when we implement the baseline method HBT on the real single-layer optical computing system; otherwise, the performance will be pretty poor. Details of our digital alignment method are in Supplement Sec. S.5.
Double-layer optical computing system: A DMD displays the binary intensity object $u_{obj}$ on the input layer. The light field is then reflected from the input layer onto the optical computing layer1 $u_{w1}$ (SLM1) and subsequently to the optical computing layer2 $u_{w2}$ (SLM2). Following this, a camera captures the modulated light present at the output layer. The length $d_{C_{1}C_{2}}$ and $d_{C_{2}O}$ for $Prop_{C_{1}C_{2}}$ and $Prop_{C_{2}O}$ in Fig. 4(b) are $215.1mm$ and $201.6mm$, respectively. More details of the optical computing system are in Supplement Sec. S.1.
3.4.2 Differentiable physics-based simulators $\hat{f}_{sys-1-layer}$ and $\hat{f}_{sys-2-layer}$.
We construct an ideal physics-based simulator $\hat{f}_{sys}$ corresponding to the aforementioned optical computing system $f_{sys}$ as the sandbox to test different training algorithms. This simulator is also used inside the design loop of SBT and HBT, which serve as the baseline methods to compare.
Since our system only includes free-space wave propagation ${\color[rgb]{1,0,0}\hat{f}_{prop}}$, wavefront modulation ${\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}$, and sensor detection ${\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}$, we build the optical computing simulator by stacking these three optical modules as building blocks. The functions of optical modules are:
${\color[rgb]{1,0,0}\hat{f}_{prop}}(u_{in},z):u_{out}=\mathcal{F}^{-1}(\mathcal{F}(u_{in})\times\mathcal{F}(h_{prop}(z))),$ (10a)
${\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}(u_{in},u_{element}):u_{out}=u_{in}*u_{element},$ (10b)
${\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}(u_{in}):I_{out}=|u_{in}|^{2},$ (10c)
where $h_{prop}(z)=\frac{e^{jkz^{\prime}}}{j\lambda z^{\prime}}e^{\frac{jk}{2z}(x^{\prime 2}+y^{\prime 2})}$ is the propagation kernel under Fresnel approximation [46] with a propagation distance $z^{\prime}$, wavelength $\lambda$ and angular wave number $k$, $u_{element}$ denotes the wavefront modulation from the programmable optical devices, $\mathcal{F}$ denotes the Fourier transform.
Figure 5: Visualization of the optical computing system’s experimental outputs and confusion matrices trained with G-MFO. (a) An input phase object digit ’$2$’ from the MNIST dataset is modulated by the optical computing layer with weight $w$ trained using G-MFO. The system correctly predicts the input as digit ‘$2$’, as the output image has the largest intensity at the region corresponding to digit ’$2$’. (b) Confusion matrix on the MNIST dataset with a training accuracy of $83.1\%$. (c) An example of a ’pullover’ from the FMNIST dataset is correctly predicted. (d) Confusion matrix on the FMNIST dataset with a training accuracy of $74.0\%$.
Simulator of a single-layer optical computing system. Based on the modules in Eq. 10, the simulator of a single-layer $\hat{f}_{sys-1-layer}=\{{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}\}$ is constructed by chaining the building blocks ${1-5}$:
$1.u_{out}={\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}(u_{laser},u_{obj}),$ (11a)
$2.u_{out}={\color[rgb]{1,0,0}\hat{f}_{prop}}(u_{out},d_{IC}),$ (11b)
$3.u_{out}={\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}(u_{out},u_{w}),$ (11c)
$4.u_{out}={\color[rgb]{1,0,0}\hat{f}_{prop}}(u_{out},d_{CO}),$ (11d)
$5.I_{cam}={\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}(u_{out}).$ (11e)
Simulator of a double-layer optical computing system is constructed similarly by stacking layers as $\hat{f}_{sys-2-layer}=\{{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}\}$:
4 Results
We aim to answer the following research questions in this section:
  • How does G-MFO perform in the numerical simulations and experiments?
  • What are the advantages and limitations of G-MFO?
  • How is G-MFO’s applicability on the marker-free cell classification task?
Thus, we numerically and experimentally evaluate the general performance of our G-MFO method on the open-source MNIST and FMNIST datasets (Subsec. 4.1). We then illustrate G-MFO’s advantage of memory- and computation-efficient training in Subsec. 4.2. Moreover, we analyze the limitation of our method in Subsec 4.3. Lastly, we demonstrate the G-MFO method on a novel application of marker-free classifying white blood cells in Subsec. 4.4.
We compare our method with the following baselines:
  • Ideal: The in silico simulation results without introducing any artificial misalignment during the simulation. This corresponds to the best achievable result for an optical computing system.
  • SBT: Simulator-based training with digitally aligned simulator.
  • HBT: The hybrid training strategies with the digitally aligned model. A detailed description of our implementation of the HBT method is in Supplement Sec. S.6.
Training details of all simulations and experiments are in Supplement Sec. S.3.
4.1 General performance evaluation on the MNIST and FMNIST dataset
4.1.1 Simulation results on the two-layer optical computing system $\hat{f}_{sys-2-layer}$.
We evaluate our method’s accuracy by conducting performance tests on a two-layer optical computing simulator $\hat{f}_{sys-2-layer}$ utilizing two classical image classification datasets: MNIST [47] and FMNIST [48]. Each optical computing layer contains $128\times 128=16,384$ trainable parameter, and the system contains $128\times 128\times 2=32,768$ parameters. Furthermore, to assess the robustness of the G-MFO method against system misalignment, we intentionally introduce a slight misalignment on the positive x’-axis direction, amounting to $40\mu m$ (equivalent to 5 pixels on SLM1) on the optical computing layer1, and $18.7\mu m$ (corresponding to 5 pixels on SLM2), on the optical computing layer2. We use $10,000$ data to train, $10,000$ data to validate, and $10,000$ data to test.
The simulation outcomes for the MNIST and FMNIST datasets are in Table 2. The ideal testing accuracy reaches $89.5\%$ on MNIST and $80.5\%$ on FMNIST. However, the presence of misalignment reduces the accuracy to $78.4\%$ and $59.8\%$ using HBT, a decrease of approximately $10\%$ and $20\%$. In contrast, the G-MFO method hits an accuracy of $87.0\%$ and $74.1\%$, effectively mitigating the detrimental impact of system misalignment.
MNIST FMNIST
Train Val Test Train Val Test
Ideal $90.0\%$ $89.3\%$ $89.5\%$ $82.2\%$ $81.2\%$ $80.5\%$
HBT $78.0\%$ $78.7\%$ $78.4\%$ $61.6\%$ $61.4\%$ $59.8\%$
G-MFO (Ours) 86.2% 86.2% 87.0% 75.4% 75.4% 74.1%
TABLE II: Numerical performance comparison on MNIST and FMNIST datasets on two-layer diffractive optical computing system.
4.1.2 G-MFO outperforms hybrid training (HBT) experimentally on a single-layer optical computing system $f_{sys-1-layer}$.
We conduct experiments on a real single-layer optical computing system $f_{sys-1-layer}$ on the MNIST and FMNIST datasets. We include in silico SBT and in situ HBT methods utilizing digitally aligned simulators as the comparison baselines. The training, validation, and testing phases each utilize a dataset of 1,000 samples.
Tab. III quantitatively shows that our method achieves higher classification accuracy than the HBT and SBT methods on both datasets in the experiments. The SBT method performs poorly due to the reality gap between the simulator and the real system. The HBT method suffers from the bias between $\hat{f}_{sys}$ and ${f}_{sys}$ in the backward process, while the G-MFO bypasses the bias-sensitive modeling and updates gradients solely with $f_{sys}$. We further show G-MFO’s capability to fine-tune the HBT result. The last row of Tab. III shows that the unbiased G-MFO method further improves the results of HBT. Moreover, we also empirically find that fine-tuning outperforms G-MFO only. Fig. 5 visualizes some experimental outputs and confusion matrices using the G-MFO method.
MNIST FMNIST
Train Val Test Train Val Test
Ideal $92.7\%$ $84.3\%$ $82.2\%$ $85.6\%$ $79.9\%$ $76.4\%$
SBT $81.9\%$ $74.4\%$ $69.3\%$ $68.3\%$ $64.1\%$ $60.9\%$
HBT $81.9\%$ $75.5\%$ $72.8\%$ $68.3\%$ $68.7\%$ $65.8\%$
G-MFO (Ours) 83.1% 77.8% 73.6% 74.0% 71.1% 70.4%
HBT+G-MFO (Ours) 87.0% 80.2% 77.7% 75.0% 71.7% 70.2%
TABLE III: Experimental performance comparison on MNIST and FMNIST datasets on the single-layer optical computing system. Results of the ideal mode are from the simulator, whose parameters are determined from experiments while we impose no misalignment on the simulator. The lower four rows are experimental results from SBT, HBT, G-MFO, and HBT+G-MFO (G-MFO fine-tuning upon the result from HBT). Our method outperforms the SBT and HBT methods on the MNIST and FMNIST datasets in experiments.
4.1.3 Experimental results on the two-layer optical computing system $f_{sys-2-layer}$.
We also perform experiments on training a real two-layer optical computing system $f_{sys-2-layer}$ on the MNIST dataset. For this experiment, we utilize a dataset comprising 10,000 instances for training, another 10,000 for validation, and a separate set of 10,000 for testing. The training, validation and testing accuracy are $80.43\%$, $81.1\%$, and $80.3\%$, respectively.
When compared with the simulated G-MFO outcomes presented in Table. II, the experimental accuracy is lower. We hypothesize that this discrepancy is attributed, in part, to mechanical disturbances encountered during the multi-day span of the experiment. Moreover, the training accuracy from two-layer system is lower than that of single-layer system while the validation and testing accuracy is higher. This is because we use more samples when training two-layer system, which alleviates overfitting.
Figure 6: All optical marker-free cell classification enabled by G-MFO training. (a) Our trained optical computing system classifies four WBC subtypes, including B cell (red), T cell (green), monocyte (blue), and granulocyte (purple), in a marker-free manner. The system uses the (i) marker-free phase information of the WBCs to (ii) perform optical classification. The output plane of the system has different regions corresponding to different classes. We (iii) measure the intensity values in these regions and (iv) choose the class with the highest intensity as the prediction. (b) shows that non-morphological features, including size and dry mass, cannot divide the four subtypes well. In contrast, (c)(i) shows the optical computing system trained with our G-MFO method can classify the WBC subtypes with a training/validation/testing accuracy of $72.1\%/73.3\%/73.8\%$, which is higher than that of the HBT method. (c)(ii) shows the confusion matrix of G-MFO results on the training set. (d) The optical computing system produces its output at the speed of light ($1.4\>ns$). This is faster than the electronic computing with ResNet10 [49, 50], which has an inference time of $1.7\>ms$.
4.2 Advantage: memory- and computation-efficient training enabled by G-MFO.
Figure 7: Advantage: G-MFO saves in silico computing resources during training. We compare the GPU (a) memory and (b) time consumption of HBT and G-MFO methods for training one batch of inputs with batch sizes of $B=4$ and $B=32$. G-MFO is more efficient than HBT during the in situ training process in memory and time. Moreover, G-MFO’s resource consumption does not increase as the batch size increases, unlike HBT.
In addition to the predicting accuracy discussed in the previous subsection, our G-MFO method has an advantage over other training algorithms regarding GPU time and memory efficiency. The SBT and HBT methods compute $\hat{f}_{sys}(x,w)$ and $\nabla_{w}\hat{f}_{sys}$ (Fig. 1b) for each input $x$ in silico, which requires a lot of in silico computation resources. In contrast, our G-MFO method executes the calculation of $f_{sys}(x,w)$ in the light-speed real optical computing system, but not in the computation- and memory-heavy simulator $\hat{f}_{sys}(x,w)$. The only step of our G-MFO method that consumes in silico computational resources is the one described in Eq. 9, where we calculate the score gradient. The in silico computational resources consumption in this step is low because it only scales with the dimension of $w$ and has no relation to the complexity of the system’s light transport $\hat{f}_{sys}$.
We compare our G-MFO method with HBT in GPU memory and time usage during experiments on a single-layer optical computing system in Fig. 7. Our G-MFO method requires far less GPU time and memory than the HBT method during the training.
4.3 Limitation: G-MFO exhibits the curse of dimensionality.
Our method is not without its limits. Our G-MFO training relies on Monte Carlo integration. It thus inherits the curse of dimensionality from the Monte Carlo integration [51]. That is, the number of samples $M$ needed to estimate the integration in Eq. 7 with a given level of accuracy grows exponentially concerning the $H$, the number of trainable parameters (i.e., dimensionality) of the function. This is also discussed and alleviated with variance reduction in the previous Sec. 3C. However, the G-MFO strategy presented in this paper is still sample-inefficient, though unbiased and memory-efficient. We need to either limit the number of trainable parameters $H$, or sample a large number of varied optical computing weights $\{w\}^{M}_{j=1}$ from distribution $p(w|\theta)$ in every iteration to make G-MFO’s gradient less noisy. The former limits our method’s design DOF, while the latter requires more executions on the real system, which prolongs the training time.
Figure 8: Limitation: MFO exhibits the curse of dimensionality. (a) When the sampling size of optical computing weights $M$ rises from $322^{5}=$ 32 to $1282^{7}=$ 128, the training accuracy increases by $2.2\%$. However, increasing $M$ from $1282^{7}=$ 128 to $2562^{9}=$ 256 only improves $0.5\%$. Here, we fix the $H$ as $128^{2}$. (b) G-MFO achieves the highest accuracy when $H=128^{2}$ under a sampling size $M=128$ while further enlarging $H$ fails the result.
We quantitatively investigate the influence of this limitation in Fig. 8 through simulations on a one-layer optical computing system $\hat{f}_{sys-1-layer}$, employing $1,000$ training samples from the MNIST dataset. The figure demonstrates how $M$ and $H$ impact the training performance of G-MFO. Shown in Fig. 8(a), G-MFO requires a $M>=128$ to achieve a training accuracy of $>90\%$ given $H=128^{2}$. Moreover, in Fig. 8(b), our G-MFO method fails when increasing $H$ beyond $128^{2}$ while keeping the sampling size $M$ fixed to $128$.
4.4 Application: all-optical classification on marker-free cellular dataset.
For the first time, we demonstrate the capability of an optical computing system for marker-free cell analysis, trained by our G-MFO algorithm (Fig. 6). We work on the white blood cells (WBC), whose abnormal subtype percentages indicate the immune system’s malfunction or infectious disease [52, 53]. We include details of the WBC phase map dataset in Supplement Sec. S.2. Previously, researchers used machine learning methods to classify WBC subtypes, including monocyte, granulocyte, B cell, and T cell, by their morphology in a marker-free way [54, 49]. However, the analysis process is computationally heavy and time-consuming.
Here, we accelerate the marker-free cell analysis process via computing with light. Our G-MFO method strikes a training/validation/testing classification accuracy of $72.1\%/73.3\%/73.8\%$ when classifying $4$ types of WBC using a one-layer optical computing system, exceeding that of the HBT method (Fig. 6c). Furthermore, Fig. 6d shows that the inference enabled by optical computing is almost instantaneous ($\frac{d_{IO}+d_{OC}}{c}=1.4\>ns$, where $c$ is the speed of light), compared to the $1.7\>ms$ of ResNet10, the electronic machine learning model used in [49]. We need $1$ more milliseconds of in silico computing of region intensities corresponding to different classes to obtain the prediction. Such a step can be skipped if we use single-photon avalanche diode (SPAD) [55] point detectors to count the corresponding regions’ cumulative signals. Though for now, the performance of our single-layer linear optical computing system is not on par with the electronic neural network, which hits a testing classification accuracy of $90.5\%$ [49], the potential of having ultra-high inference speed and $>70\%$ classification accuracy here point out an exciting direction on further increasing the complexity of our optical computing system to improve the absolute classification accuracy on classifying cells.
5 Discussion
5.1 Future directions.
5.1.1 Further apply G-MFO to more complex optical computing systems.
Exploring the scalability of the optical computing system to more complex optical structures, along with integrating more layers and non-linear activation functions, could potentially enhance absolute performance.
5.1.2 Further improve the performance of G-MFO.
Future research could also consider employing more advanced techniques related to Monte Carlo integration to reduce the training variance discussed in the previous Subsec. 4.3, which we anticipate could substantially broaden the viable search space, thus further empowering the G-MFO approach. These include using more advanced sampling strategies [56] or integrating G-MFO with the SBT methods [57] or adding critic function [58]. The latter two methods trade off the introduced bias and sampling variance.
5.1.3 Expanding the application of G-MFO to additional computational optics tasks.
The concept of G-MFO presents promising avenues for application in other areas of computational optics, such as computer-generated holography [59] and lens design [60]. The inherent model-free and resource-efficient characteristics of G-MFO position it as a viable alternative to prevalent model-based methods [61, 60]. Future research could focus on leveraging G-MFO to these domains, potentially enhancing computational efficiency and performance.
5.2 Conclusion
To conclude, our study underscores the effectiveness of a model-free strategy in training optical computing systems in situ, manifesting considerable potential in computational efficiency and reducing the simulation-to-reality performance gap. Although the study does not focus entirely on absolute image classification accuracy as it is based on a simple single or double diffractive optical computing system without non-linearity, it shows relative improvements compared to the existing training strategies, indicating that our strategy is a potentially valuable approach. The model-agnostic nature of our technique may become even more beneficial when implemented in intricate optical systems, representing a robust and versatile alternative to current strategies. It promises a strong foundation for exploring and practically implementing optical computing in real-world applications such as high-speed marker-free cell analysis.
Funding
Hong Kong General Research Fund (14209521); Hong Kong Innovation and Technology Fund (ITS/178/20FP & ITS/148/20); Croucher Foundation (CM/CT/CF/CIA/0688/19ay).
Acknowledgments
We thank Cheng Zheng for the discussions in the early stages of the work.
Author Contributions. G.Z. conceived the project, derived the formulation, and built the backbone code and system. X.S. helped with the code writing and system setup and collected the simulation and experiment results. R.Z. supervised the project. G.Z. and X.S. wrote the manuscript with comments and edits from R.Z..
Supplemental document
See Supplement for supporting content.